# Incorporating Judgment Prediction into Legal Case Retrieval via Law-aware Generative Retrieval

### v1

@@ 3.3. Ablation Studies
Firstly, we test the performance of Gear’s variants by removing a certain component. Following DSI (Tay et al., 2022), we replace rationale extraction and law-aware hierarchical identifiers to the Direct indexing (the first 512 tokens) and $k$-means identifiers. As shown in the bottom part of Table 2, we observe that: (1) Removing modules $\mathcal{L}_{c}$, $f_{R}$, and $f_{T}$ individually results in a performance degradation of 25.51%, 23.83%, and 31.16% on ELAM, 8.50%, 25.00%, and 26.13% on LeCaRDv2. These results verify the effectiveness of all three modules. (2) It is worth noting that removing our law-aware hierarchical identifiers results in the most significant performance decrease. It is because we inject law knowledge into each digit of the identifier, aligning them with the hierarchical structure of laws. By simulating the retrieval to mirror the legal decision-making process, we enable Gear to effectively learn the association between query cases and candidate cases. These results demonstrate the effectiveness of our identifiers and highlight the importance of introducing the law knowledge to legal case retrieval.
To further validate the efficacy of the components of Gear, we combine the three components of Gear with DSI individually and test the retrieval performance. We conduct 5 repeated experiments and present the mean values of MRR, with error bars representing the 95% confidence interval. As illustrated in Figure 4, incorporating our components of Gear into DSI consistently results in performance improvements. Specifically, DSI equipped with rationale extraction achieves an improvement of over 60% on ELAM, and DSI equipped with law-aware identifiers demonstrated a 42% improvement on LeCaRDv2 dataset. However, DSI equipped with revision loss does not exhibit a substantial improvement. Because DSI utilizes $k$-means identifiers and the length of these identifiers may vary due to differences in cluster sizes. In such cases, our revision loss cannot accurately measure the differences between predictions and labels, resulting in a slight performance improvement.
(a) $coverage_{\text{freq}}$ on LeCaRDv2
(b) $coverage_{\text{top1}}$ on LeCaRDv2
(c) $coverage$ on ELAM
@@ 3.4. Quality of Judgment Prediction
To answer RQ3, we conduct experiments to evaluate the judgment prediction accuracy of retrieved cases using the proposed metric $coverage$. Considering that in LeCaRDv2, the ground-truth judgment label of query cases is not provided, we use $coverage_{\text{freq}}$ and $coverage_{\text{top1}}$ to create pseudo labels for them. We run experiments ten times and present the results in Figure 5, reporting the average score with the shaded area indicating the 95% confidence interval. From the plot, we observe that: (1) in the single-charge scenario as shown in Figure 5(c) , Gear achieves a remarkably high $coverage$ score (over 0.95) with just 1 case retrieved. In the multi-charge scenario as shown in Figure 5(a) and Figure 5(b), Gear retrieve about 3 cases to encompass approximately 85% of the charges. These results demonstrate that Gear has considerable ability in charge prediction. This capability stems from our integration of judgment prediction into case retrieval, unifying the predictions for both tasks in a traversal on the law structure constraint tree. In other words, we not only require the model to retrieve relevant cases but also ensure that the retrieved cases share the same applicable charges as the query case. This incorporation brings the judgment prediction capability to Gear. (2) Gear demonstrates a significantly coverage improvement compared to DSI and DSI-QG on both datasets, especially when a limited number of cases are retrieved, such as 1 or 3. This mainly attributes to the specialized design of Gear for judgment prediction tasks including the law-aware hierarchical identifiers and the revision loss, which explicitly enhances the accuracy of judgment predictions. These experimental results verify Gear is capable of performing competitive legal judgment predictions.
3.5. Quality of Extracted Rationales
@@
To answer RQ4, we conduct experiments and evaluate the rationales extracted by Gear in comparison to those generated by prevalent query generation methods featured in existing generative retrieval studies. We consider the Direct indexing (Tay et al., 2022), which means using the first 512 tokens as the query; Doc2query (Zhuang et al., 2022; Wang et al., 2022), which uses a language model to generate pseudo text as the model input; and ChatGPT, which demonstrates to provide text summaries that are on par with human-level quality (Zhang et al., 2023b; Bang et al., 2023; Liu et al., 2023b; Wang et al., 2023). First, we follow the practice (Zhao and Vydiswaran, 2021; Yu et al., 2022) and conduct the human evaluation to assess the quality of extracted rationales. We randomly selected 50 examples from both ELAM and LeCaRDv2 and ask two annotators with a legal background to determine whether the extracted rationales (generated queries) are sufficient to ascertain the applicable charges for the original cases, indicating that the extracted sentences comprehensively encompass the primary information from the original cases. Each annotator is provided with the rationales (generated queries) and the corresponding original case and is asked to mark the question as 1 if they agreed with it and 0 otherwise. As shown in Table 3, the results illustrate rationales extracted by Gear consistently outperform both Direct indexing and Doc2query on two datasets, achieving performance comparable to that of ChatGPT. These results confirm the rationales extracted by Gear is informative for both tasks.
@@
Table 3. Time overhead and human evaluation of rationale extraction over 50 randomly sampled data from ELAM and LeCaRDv2 by two annotators with the inter-rater agreement of 0.96. “Time”(ms) denotes the extraction time per case.
@@
Table 4. Comparison in average inference wall time (95% confidence interval) per query on a single Nvidia A6000.
@@
Then, we also compare the time overhead of Gear for extracting rationales and baselines for generating queries. Please note that Direct indexing, as it directly takes the first 512 tokens as the query, has negligible time overhead. Besides, the runtime of ChatGPT API may be affected by network latency, we thus conduct 50 tests and report the average time overhead and its 95% confidence interval as shown in Table 3. From the results, we can see that Gear achieves performance similar to ChatGPT but with far less time consumption. Compared to Doc2query, Gear also exhibits a significant advantage in terms of time overhead.
To gain a more intuitive understanding of the differences among the texts generated by various methods, we conduct a case study in Appendix B.1.
4. Inference Time and Convergence
@@
To answer RQ5, we record the inference wall time (in milliseconds) per query of baseline methods and Gear. This comparison is conducted on a single Nvidia A6000 across two datasets. From the results shown in Table 4, we observe that Gear demonstrates superior efficiency, with an inference time of 59.184 ms (±8.737) for the ELAM dataset, and 20.391 ms (±4.064) for LeCaRDv2. Specifically, when compared to Ultron, Gear achieves a remarkable 46.48% reduction in inference time for the ELAM dataset and a 36.03% reduction for LeCaRDv2. This is mainly because Ultron uses product quantization to create the identifier for each case. For all case embeddings, Ultron first divides embedding space into several groups and then performs $k$-means clustering on each group. It usually leads to excessively long identifiers. In the case of DSI and DSI-QG, Gear exhibits a substantial 19.01% and 18.59% improvement for ELAM and a 22.64% and 18.59% improvement for LeCaRDv2. The reason for this result is that in DSI and DSI-QG the tree constructed by $k$-means may be unbalanced, meaning that the lengths of case identifiers are unequal. Some case identifiers may be longer, which impairs the inference performance. These findings indicate the remarkable efficiency of Gear.
@@
For further confirming Gear’s efficiency, we plot the testing curves for DSI, DSI-QG, and our Gear with the $x$-axis denoting the number of epochs, the $y$-axis denoting the MRR score, and shaded area indicating the 95% confidence interval. As shown in Figure 6, we observe that Gear not only outperforms DSI and DSI-QG significantly in terms of performance but also exhibits superior convergence. Gear achieves near-optimal MRR performance by 5 epochs, whereas DSI-QG and DSI converge at 6 and 8 epochs, respectively. The results verify the efficiency of Gear.
5. Related Work
5.1. Legal Case Retrieval
Legal case retrieval is a long-standing research topic in the field of information retrieval. In the early exploration, researchers (Moens, 2001; Zeng et al., 2005; Klein et al., 2006; Saravanan et al., 2009; Bench-Capon et al., 2012) made efforts to inject the legal knowledge to the case retrieval through the decomposition of legal issues and involving the ontology. In recent years, deep learning techniques have demonstrated their effectiveness in exploring the semantics of legal case retrieval. One representative line of works focused on the network-based precedents methods tailored for the common law systems. For example, Minocha et al. (2015) leveraged the Precedent Citation Network (PCNet) to predict the relevance based on whether the sets of precedent citations occur in the same cluster. Bhattacharya et al. (2020) proposed Hier-SPCNet to capture all domain information inherent in laws and precedents. The other line of works judged the relevance between the query case and candidate case according to their text-level similarity. One such method is BERT-PLI (Shao et al., 2020). It divided the legal case texts into several paragraphs and used BERT (Devlin et al., 2018) to obtain the similarity between the paragraphs, achieving promising ranking performance. Lawformer (Xiao et al., 2021b) is another text-based method. It used millions of Chinese criminal and civil case documents to pre-train a longformer (Beltagy et al., 2020) model and achieved impressive performance across legal tasks. To further estimate the relevance score in the legal domain, Li et al. (2023) made full use of the structure relationships of legal case documents for pre-training and achieved state-of-the-art performances.
Although these studies have made efforts to improve the retrieval performance, they ignored the charge-level consistency between the query case and the retrieved case, resulting in sub-optimal and unreliable results.
@@ 5.1. Legal Case Retrieval
5.2. Generative Retrieval
Generative retrieval has recently emerged as a promising direction for document retrieval, enjoying merits like being end-to-end differentiable. Various methods have been introduced to generate semantic document identifiers. For example,  Cao et al. (2021) introduced a related sequence-to-sequence system to conduct entity retrieval. They first represented documents as unique names that are composed of entity names. Then they used a auto-regressive generation model to generate the unique names of these entities based on contextual information.  Tay et al. (2022) proposed the differentiable search index (DSI) paradigm, which is an auto-regressive generation model to perform ad-hoc retrieval tasks. The input of the model was a natural language query and the model regressively generated documents’ identifier strings that are relevant to the given query.  Wang et al. (2022) proposed a novel method NCI, which used a tailored prefix-aware weight-adaptive decoder to optimize the recall and ranking performance. NCI also used constrained beam search to make sure all the generated identifiers were valid. Ultron (Zhou et al., 2022) leveraged document titles and substrings as identifiers to enrich the semantic information of identifiers. To mitigate data distribution mismatch that occurs between the indexing and the retrieval phases,  Zhuang et al. (2022) proposed DSI-QG, which adopted a query generation model with a cross-encoder ranker to generate and select a set of relevant queries.
Existing studies focused on the general documents, lacking specific designs for legal cases, particularly the incorporation of law knowledge. As a result, it fails to deliver satisfactory performance.
@@ 6. Conclusion
In this study, we introduce Gear, a novel generative legal case retrieval method that integrates judgment prediction into the retrieval process. Gear explicitly exploits the law knowledge and extracts rationales from legal cases, ensuring a shared and informative representation for both tasks. Grounded with the inherent hierarchy of laws, Gear constructs a law structure constraint tree and assigns the law-aware semantic identifier to each case. These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which respectively correspond to thses two tasks. With the help of the proposed revision loss, Gear jointly minimizes the discrepancy between the identifiers of predicted and labeled charges/ retrieved cases, improving the accuracy and consistency for both tasks. Extensive experiments on two datasets demonstrate that Gear consistently outperforms state-of-the-art methods in legal case retrieval while maintaining competitive judgment prediction performance.
Appendix A REPRODUCIBILITY
A.1. Datasets
We conduct experiments on ELAM (Yu et al., 2022) and LeCaRDv2 (Ma et al., 2021) for legal case retrieval in this work.
ELAM (Yu et al., 2022) [^5] https://github.com/ruc-wjyu/OPT-Match. is a Chinese legal case retrieval dataset, focusing on criminal cases. ELAM has corresponding labels for both legal case retrieval and judgment prediction, which is suitable for our goal. We exclude those cases with multiple applicable charges to consider the retrieval and judgment prediction performance in a single charge scenario. The resulting candidate pool size of ELAM is 1332. Other data preprocessing is aligned to (Li et al., 2023).
LeCaRDv2666https://github.com/THUIR/LeCaRDv2. is the official updated version of LeCaRD  (Ma et al., 2021). In this dataset, the relevance labels are divided into four levels, ranging from 3 to 0, indicating a gradual decrease in relevance. We follow the data preprocessing approach of (Li et al., 2023), with the exception of increasing the candidate pool size from 100 to 1390 to further validate the effectiveness of baselines and our model. In LeCaRDv2, we consider both single and multiple charges, where each case corresponds 1.5 charges on average. Although LeCaRDv2 does not provide the judgment labels, we instead create pseudo labels using the proposed metric $coverage$ in Section 3.1.
A.2. Baselines
We consider three types of baselines in this study.
(1) Term-based methods:
  • Query Likelihood (QL) (Zhai et al., 2008) is a probabilistic language modeling approach used to estimate the relevance of documents to a given query. In QL, the likelihood of a document generating the words in a query is calculated, treating both the query and the document as random variables. This approach is based on the assumption that relevant documents are more likely to generate the query terms.
  • BM25 (Robertson et al., 2009) is a probabilistic information retrieval model widely used in the field of text retrieval. BM25 takes into account both term frequency and document length normalization, making it a robust and effective ranking model.
(2) Dense retrieval methods:
  • ANCE (Xiong et al., 2020) is a contrastive learning based model for ad-hoc dense retrieval. It selects hard training negatives from the entire corpus to reduce the stochastic gradient variance, and improve retrieval performance.
  • BERT+ANN (Devlin et al., 2018; Chalkidis et al., 2020) is a strong baseline in ad-hoc retrieval tasks that uses BERT to encode documents and then apply Approximate Nearest Neighbor (ANN) search algorithms to retrieve relevant document. In this work, we adopt the BERT pre-trained on a large Chinese legal corpus [^7] https://github.com/thunlp/OpenCLaP..
  • Lawformer+ANN (Xiao et al., 2021a): similar to BERT+ANN, we adopt Lawformer, which is a longformer (Beltagy et al., 2020) backbone and pre-trained on large legal case corpus, to encode legal cases.
  • SAILER (Li et al., 2023) is a structure-aware pre-training method for legal case retrieval. It adopts an asymmetric encoder-decoder architecture to integrate structures of legal case document information into dense vectors. SAILER is the state-of-the-art legal case retrieval method.
(3) Generative retrieval methods:
  • DSI (Tay et al., 2022) is a new paradigm for document retrieval tasks. It utilizes a Transformer-based encoder-decoder model to map queries directly to relevant identifiers. DSI achieves the end-to-end retrieval.
  • NCI (Wang et al., 2022) improves DSI in terms of using constrained beam search and prefix-aware weight-adaptive decoder.
  • DSI-QG (Zhuang et al., 2022) design a query generation process on the top of DSI, which can mitigate data distribution mismatches present between the indexing and the retrieval phases.
  • Ultron (Zhou et al., 2022) improves DSI through adopting product quantization to create semantic-based identifier and using URL to create term-based identifier.
A.3. Implementation Details
We implement baseline methods following the suggestions in the original papers. Specifically, (1) For classical term-based baselines, we use the pyserini [^8] https://github.com/castorini/pyserini. and genism [^9] https://github.com/RaRe-Technologies/gensim. toolkits with the default parameters. (2) For dense retrieval baselines, we use the open-sourced checkpoint [^10] https://github.com/thunlp. to initialize model parameters of BERT and Lawformer and use faiss [^11] https://github.com/facebookresearch/faiss.toolkit to implement ANN algorithms. The batch size is set to 16, the max length of input case is set to 512 for BERT and 1024 for Lawformer. We tune these models with in-batch contrastive loss. (3) For generative retrieval baselines, we directly use their official open-source implementations with the Chinese pre-trained T5 “Randeng”  [^12] https://huggingface.co/IDEA-CCNL/Randeng-T5-77M-MultiTask-Chinese. as backbone. We use beam search to retrieve relevant cases, where the beam size is set to 30.
We keep the backbone model and beam size same as baselines and set the max input length of Gear to 512, the rest hyperparameters are tuned as follows: the batch size is set to 2; the learning rate is tuned from [1e-5,1e-4] with step size 2e-5; in the rationale extraction module, $k_{1}$, $k_{2}$, $k_{3}$ are respectively set to 2, 5, 15 for ELAM and 10, 20, 15 for LeCaRDv2 ; $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are respectively set to 10.0, 1.0, 0.1; in the law structure constraint tree module, the height of the tree (the length of the hierarchical identifier) $L$ is set to 4; in the training, $\mu$ is set to 1; $\gamma_{h}$ is tuned from $\{0.01,0.1\}$; $\gamma_{h}$ is set to 1; $\lambda_{w}$ is tuned from $\{1,10,100\}$; $\lambda_{l}$ is tuned from [1e-4,1e-2] with step size 5e-4. Hyperparameters of Gear are tuned using grid search with Adam(Kingma and Ba, 2014). ALL experiments are conducted on a single Nvidia A6000.
Figure 7. Comparison between the rationale extracted by Gear and the queries generated by Direct indexing, Doc2query, and ChatGPT. Underlined snippets denotes the key circumstances or key elements for the case.
Figure 8. A case study of prompting ChatGPT to obatin the key elements and key circumstances for a charge.
Appendix B Supplementary Experiments
B.1. Case Study in Rationale Extraction
In this section, we employ a case study to compare the performance of Gear and baseline methods in rationale extraction (query generation). Figure 7 showcases the comparison with all results translated from Chinese. We can find that (1) queries generated by Direct indexing and Doc2query fail to articulate the key circumstances and key elements of the original case, which are crucial factors for evaluating the relevance in legal case retrieval. This is because they lack law knowledge and struggle to handle the extremely lengthy documents. (2) The quality of both rationales extraction and ChatGPT summary is considerable as the underlined snippets highlighted. While both convey a substantial amount of key information, the rationales extracted by Gear are more sufficient and organized, which is easier for model learning. It contributes to our multi-level extraction strategy where we respectively extract keywords and sentences according to the definition of laws. The above results validate the effectiveness of the rationale extraction of Gear.
B.2. Building Corpus by Prompting ChatGPT
To improve the effectiveness of rationale extraction, we prompt chatGPT and extend the corpus using chatGPT’s feedback. Our prompt is shown in Figure 8. We would like ChatGPT play the role of a legal expert targeting on the Criminal Law and ask ChatGPT output the key circumstances and key elements for each charge. From the case, we found that chatGPT can provide multiple perspectives on a charge, including perpetrator, victim, behavior, demands, etc. These annotations serve as supplementary clarifications to the Criminal Law definitions and help Gear accurately select the informative rationales from legal cases.

### v2

@@ 4.3. Ablation Studies
Firstly, we test the performance of GEAR’s variants by removing a certain component. Following DSI (Tay et al., 2022), we replace rationale extraction and law-aware hierarchical IDs with Direct indexing (using the first 512 tokens of document as queries for language models) and $k$-means IDs. As shown in the bottom part of Table 2, we observe that: (1) Removing modules $\mathcal{L}_{c}$, $f_{R}$, and $f_{T}$ individually results in a performance degradation of 25.51%, 23.83%, and 31.16% on ELAM, 8.50%, 25.00%, and 26.13% on LeCaRDv2. These results verify the effectiveness of all three modules. (2) It is worth noting that removing our law-aware hierarchical IDs results in the most significant performance decrease. It is because we inject law knowledge into each digit of the ID, aligning them with the hierarchical structure of laws. By simulating the retrieval to mirror the legal decision-making process, we enable GEAR to effectively learn the association between query cases and candidate cases. These results demonstrate the effectiveness of our IDs and highlight the importance of introducing structural and semantic knowledge in law to legal document retrieval.
To further validate the efficacy of the components of GEAR, we combine the three components of GEAR with DSI individually and test its retrieval performance. We conduct 5 repeated experiments and present the mean values of MRR, with error bars representing the 95% confidence interval. As illustrated in Figure 3, integrating the components of GEAR into DSI consistently results in performance improvements. Specifically, DSI equipped with rationale extraction achieves an improvement of over 60% on ELAM, and DSI equipped with law-aware IDs demonstrated a 42% improvement on LeCaRDv2 dataset. However, DSI equipped with revision loss does not exhibit a substantial improvement. Because DSI utilizes $k$-means IDs and the length of these IDs may vary due to differences in cluster sizes. In such cases, our revision loss cannot accurately measure the differences between predictions and labels, resulting in a slight performance improvement.
@@ 4.4. Quality of Judgment Prediction
To answer RQ3, we conduct experiments to evaluate the judgment prediction accuracy of retrieved cases using the proposed metric $coverage@k$. Please note that we consider the applicable charges as the judgment results.
@@
We run experiments ten times and present the results in Figure 4, reporting the average score with the shaded area indicating the 95% confidence interval. From the plot, we observe that: (1) in the single-charge scenario (ELAM) as shown in Figure 4(a) , GEAR achieves a remarkably high $coverage$ score (over 0.95) with just 1 case retrieved. In the multi-charge scenario (LeCaRDv2) as shown in Figure 4(b), GEAR retrieves about 3 cases to encompass approximately 85% of the charges. These results demonstrate that GEAR has considerable ability in charge prediction. This capability stems from our integration of judgment prediction into case retrieval, unifying the predictions for both tasks in a traversal on the law structure constraint tree. (2) GEAR demonstrates a significantly coverage improvement compared to DSI and DSI-QG on both datasets, especially when a limited number of cases are retrieved, such as 1 or 3. This mainly attributes to the specialized design of GEAR for judgment prediction tasks including the law-aware hierarchical IDs and the revision loss, which explicitly enhances the accuracy of judgment predictions. These results verify GEAR is capable of performing competitive legal judgment predictions.
4.5. Effectiveness and Efficiency of Rationale Extraction
@@
To answer RQ4, we conduct experiments and evaluate the rationales extracted by GEAR in comparison to those generated by prevalent query generation methods featured in existing generative retrieval studies. We consider the following baselines: Direct indexing (Tay et al., 2022), which means using the first 512 tokens of the raw document as the query; Doc2query (Zhuang et al., 2022; Wang et al., 2022), which uses a language model to generate pseudo text as the model input; and ChatLaw (Cui et al., 2023), which demonstrates to provide legal summaries that are on par with human-level quality.
First, we follow the practice (Zhao and Vydiswaran, 2021; Yu et al., 2022b) and conduct the human evaluation to assess the quality of extracted rationales. We randomly selected 50 samples from both ELAM and LeCaRDv2 and asked two annotators (Ph.D. in Law) to determine whether the extracted rationales (generated queries) are sufficient to ascertain the applicable charges for the original cases, indicating that the extracted sentences comprehensively encompass the primary information from the original cases. Each annotator is provided with a pair of the rationales (generated queries) and the corresponding original case for each sample and is asked to mark the sample as 1 if they agreed with it and 0 otherwise. We calculate the accuracy of the sample rationales based on annotators’ evaluation. As shown in Table 3, the results illustrate rationales extracted by GEAR consistently outperform both Direct indexing and Doc2query on two datasets, achieving performance comparable to that of ChatLaw. These results confirm the rationales extracted by GEAR are informative for both tasks.
Then, we conduct an experiment to validate the time overhead of rationale extraction. Since Direct indexing takes the first 512 tokens as the query, we omit its time overhead. From the results shown in Table 4, we can see that GEAR achieves impressive accuracy on par with ChatLaw but with far less time consumption. Compared to Doc2query, GEAR also exhibits a significant advantage in terms of time overhead. These results verify the GEAR’s efficiency in terms of rationale extraction. Based on the results, the generative approaches for rationales are not advisable for legal document retrieval due to the efficiency issue.
@@
Table 3. Time overhead and human evaluation of rationale extraction over 50 random samples from ELAM and LeCaRDv2 by two annotators with the inter-rater agreement of 0.96. “Time”(ms) denotes the extraction time per sample.
@@
Table 4. Comparison in average inference wall time (95% confidence interval) per query on a single NVIDIA RTX A6000.
@@
Table 5. Performance comparisons of our approach and the baselines on BSARD. “R@K” is short for “Recall@K”. indicates that improvements are significant based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
Models R@5 R@10 R@20 R@30 MRR
QL (Zhai et al., 2008) 0.1752 0.2128 0.2698 0.2923 0.1770
BM25 (Robertson et al., 2009) 0.1815 0.2381 0.2871 0.3056 0.1786
G-DSR (Louis et al., 2023) 0.1636 0.3426 0.5081 0.6720 0.3012
DSI (Tay et al., 2022) 0.4578 0.5536 0.6315 0.6562 0.3955
GEAR 0.5297 0.6164 0.7031 0.7170 0.4081
4.6. Inference Time and Convergence
@@
To answer RQ5, we record the inference wall time (in milliseconds) per query of baseline methods and GEAR. This comparison is conducted on a single NVIDIA RTX A6000 across two datasets. From the results shown in Table 4, we observe that GEAR demonstrates superior efficiency, with an inference time of 59.184 ms (±8.737) for the ELAM dataset, and 20.391 ms (±4.064) for LeCaRDv2. Specifically, when compared to Ultron, GEAR achieves a remarkable 46.48% reduction in inference time for the ELAM dataset and a 36.03% reduction for LeCaRDv2. This is mainly because Ultron uses product quantization to create the IDs for documents. For all document embeddings, Ultron first divides embedding space into several groups and then performs $k$-means clustering on each group. It usually leads to excessively long IDs. In the case of DSI and DSI-QG, GEAR exhibits a substantial 19.01% and 18.59% improvement for ELAM and a 22.64% and 18.59% improvement for LeCaRDv2. In DSI and DSI-QG, the tree constructed by $k$-means may be unbalanced, meaning that the lengths of case IDs are unequal. Some case IDs may be longer, which impairs the inference performance.
@@
For further confirming GEAR’s efficiency, we plot the testing curves for DSI, DSI-QG, and our GEAR with the $x$-axis denoting the number of epochs, the $y$-axis denoting the MRR score, and the shaded area indicating the 95% confidence interval. As illustrated in Figure 5, we observe that GEAR not only outperforms DSI and DSI-QG significantly in terms of performance but also exhibits superior convergence. GEAR achieves near-optimal MRR performance by 6 epochs, whereas DSI-QG and DSI converge at 8 and 10 epochs, respectively. The results verify the efficiency of GEAR.
@@
4.7. Robustness across Languages and Domains
To answer RQ6, we compared the performance of baselines and GEAR on BSARD, a French SAR dataset. Since BSARD has a relatively short query length, we omit the rationale extraction part and take the raw legal questions as queries for GEAR. In terms of the ID, we follow the structure of the Belgian code provided in the data and assign the hierarchical semantic to each statutory article. From the results shown in Table 5, we have two observations: (1) with a small number of documents to retrieve, generative retrieval methods (DSI and GEAR) exhibit significantly higher retrieval performances in SAR scenario compared to dense retrieval. We assume the reason why dense models perform poorer is that there exists a significant gap between legal questions and statutory articles. It is difficult for dense models to learn the correct association between them especially without the law knowledge injected. (2) GEAR demonstrates the best performance, exhibiting a substantial advantage over sparse and dense retrieval methods including the current state-of-the-art model G-DSR. The improvement of GEAR benefits from explicitly injecting legal knowledge into generative retrieval frameworks.
@@ 5. Conclusion
In this study, we introduce GEAR, a novel law-guided generative legal document retrieval method that explicitly integrates judgment prediction. GEAR exploits the law knowledge and extracts rationales from legal documents, ensuring a shared and informative representation for both tasks. Grounded in the inherent hierarchy of laws, GEAR constructs a law structure constraint tree and assigns the law-aware semantic ID to each document. These designs enable a unified traversal from the root, through intermediate charge nodes, to case-specific leaf nodes, which empowers GEAR to perform dual predictions for judgment and relevant documents in a single inference. With the help of the proposed revision loss, GEAR jointly minimizes the discrepancy between the IDs of predicted and labeled judgments/ retrieved documents, improving the accuracy and consistency for both tasks. Extensive experiments on two LCR datasets show the superiority of GEAR over state-of-the-art methods while maintaining competitive judgment prediction performance. Moreover, we validate its robustness across languages and domains on a French SAR dataset.
Acknowledgements.
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62376275, No. 62377044), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China.